OpenStudy (anonymous):

What is .... @ganeshie8

OpenStudy (anonymous):

What is\[5\pi/3\] radians in degrees?

ganeshie8 (ganeshie8):

radians -> degrees : multiply by 180/pi

OpenStudy (anonymous):

300?

ganeshie8 (ganeshie8):

\(\large \frac{5\pi}{3} ~radians = \frac{5 \pi}{3} \times \frac{180}{\pi}~ degrees \)

ganeshie8 (ganeshie8):

300 degrees is \(\large \color{red}{\checkmark}\)